Solve the proportion please:
5/6 = a-2/4

Answer by xcentaur(357)

5/6 = a-2/4


5(4) = 6(a-2)


20 = 6a - 12


20+12 = 6a


6a=32


a=32/6
a=16/3


Ans:a=16/3